OpenStudy (amenah8):

find the local maximum and local minimum for each if they exist. (INCLUDE THE VALUES) f(x)=(1/4)x^4 - (5/3)x^3 + 3x^2

OpenStudy (mathmale):

First, I'd like to compliment you on having enclosed those fractions in parentheses. Very appropriate, very clear. :) I assume this problem is one from calculus. The first step of applying calculus to finding the local min and the local max (if either exists) is to differentiate the given function with respect to x. Please do that, and write your derivative here: f '(x) = ?

OpenStudy (amenah8):

ok, i'll do that then get back with you!

OpenStudy (amenah8):

f prime of x = x^3 - 5x^2 + 6x

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

did you get the derivative ?

OpenStudy (misty1212):

oh yes, i see that you did set it equal to zero and solve

OpenStudy (misty1212):

you are lucky because this one factors

OpenStudy (misty1212):

you good from there or no?

OpenStudy (amenah8):

i think so. sorry, i was trying to do it based on what you said! one moment please...

OpenStudy (amenah8):

x^2 (x-5+6)=0 ??

OpenStudy (anonymous):

nope only x out front, not x^2

OpenStudy (anonymous):

\[x^2-5x+6)=0\]

OpenStudy (anonymous):

ok not that \[x(x^2-5x+6)=01\]

OpenStudy (anonymous):

the second part factors,so the zeros are easy to find

OpenStudy (amenah8):

so now we are supposed to factor what is in the parenthesis?

OpenStudy (anonymous):

yes that is easier than using the quadratic formula in this case

OpenStudy (amenah8):

but you can't factor an x out of 6? what would be the common factor?